OpenStudy (anonymous):

f(x)=eˆ(x/2)-2 ; is the zero points to the function =-1 or 2 ?????

OpenStudy (badhi):

are you asing the value \(x_0\) where\(f(x_0) = 0\) ?

OpenStudy (anonymous):

yes

OpenStudy (badhi):

so start with \(f(x)=0\) then youll get, \[e^{\frac x 2 }-2 = 0\] did you solve this??

OpenStudy (anonymous):

but i think that e^0 =1 so 1-2=-1 or what?

OpenStudy (anonymous):

yes i solve it like this

OpenStudy (badhi):

so you end up with -1 . but you should end up with 0. I mean \[e^{(\text{something} )}-2 =0 \] you chose that something as 0 but ended up with -1 but it should be 0 Rather than guessing values I recommend following steps \[\begin{align} e^{\frac x 2} -2 &=0 \\ e^{\frac x 2} &= 2 \\ \ln\left(e^{\frac x 2}\right)=\ln 2 \end{align}\] can you continue and solve for x?

OpenStudy (badhi):

yep

OpenStudy (anonymous):

is that the only zero point?

OpenStudy (badhi):

looks so to me :)